# Mean-field Potts and random-cluster dynamics from high-entropy initializations

### v1

@@ 1.3 Initializations at other temperatures and relations to simulated annealing
Another central purpose of understanding mixing times of high-entropy initializations for Markov chains is that it is closely related to understanding simulated annealing schemes. Simulated annealing, introduced in [36] (and variants like simulated tempering [41]) lowers the temperature over the run of the Markov chain, so that when run at low-temperatures, one is effectively initializing from the stationary measure at a nearby, but higher temperature (and therefore higher-entropy) distribution. This is one of the most commonly implemented modifications to vanilla MCMC to make it better able to sample from multimodal low-temperature landscapes. Simulated annealing and simulated tempering have been analyzed in the context of mean-field Potts dynamics by e.g., [1, 47] and it is known that the discontinuous phase transition at $\beta_{\mathsf{c}}$ presents a serious obstruction to fast mixing from such schemes, for instance because the high-temperature initialization is metastable when $\beta\in(\beta_{\mathsf{c}},\beta_{\mathsf{s}})$.
Still, our results lend insight into the approach by precisely classifying the set of temperatures $\beta_{0}$ for which initialization from $\mu_{\beta_{0}}$ or $\pi_{\beta_{0}}$ would be fast for the CM and Glauber dynamics. Because the properties of the initialization used in our proofs are only quantities like the size of the giant component, the sum of squares of non-giant component sizes for random-cluster and the proportions vector in the Potts Glauber case, applying them to initializations from $\mu_{\beta_{0}}$ or $\pi_{\beta_{0}}$ instead of the product measures yields the following: there exists an explicit $b_{*}(\beta,q)\geq\beta_{\mathsf{c}}$ such that
  • When $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$, the CM dynamics initialized from $\mu_{\beta_{0}}$ for $\beta_{0}<b_{*}(\beta,q)$ will mix in $O(\log n)$ steps, but will be slow from any $\beta>b_{*}(\beta,q)$.
  • When $\beta=\beta_{\mathsf{c}}$, the CM dynamics initialized from $\mu_{\beta_{0}}$ will be slow to equilibrate for every $\beta_{0}\neq\beta_{\mathsf{c}}$.
  • When $\beta\in(\beta_{\mathsf{c}},\beta_{\mathsf{s}})$, the CM dynamics initialized from $\mu_{\beta_{0}}$ mixes in $O(\log n)$ steps for every $\beta_{0}>b_{*}(\beta,q)$ but not from $\beta_{0}<b_{*}(\beta,q)$.
Similar results hold for simulated annealing for the Potts Glauber dynamics, where per item 4 of Theorem 1.2 one also gets $O(n\log n)$ mixing at low-temperature $\beta>\beta_{\mathsf{s}}$ when initialized from $\pi_{\beta_{0}}$ for any $\beta_{0}>0$.
@@ 1.4 Proof outlines
In what follows, let $f(\theta)=f_{\beta,q}(\theta)$ be the expected drift for the size of the giant component in one step of CM dynamics. When $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$, let $\theta_{*}$ be its unstable fixed point separating the two stable fixed points of 00 (the disordered phase) and $\theta_{\mathsf{r}}$ (the ordered phase); see Figure 3 and Section 2.3 for the precise definitions. For intuition, when comparing to Figure 1, the drift is roughly the derivative of the log-probability landscape: a giant component of size $\theta_{*}n$ corresponds to the saddle point separating the modes whose minima are at no giant component, and at a giant component of size $\theta_{\mathsf{r}}n$.
@@
    Escape from a local neighborhood of the unstable fixed point: We show that locally near, i.e., in an $O(\sqrt{n})$ window around the unstable fixed point $\theta_{*}n$, the evolution of the size of its giant component, though not literally a Markov process, is well-approximated by a 1-dimensional Gaussian Markov chain (so long as the dynamics maintains a certain typical structure).
    The 1-dimensional Gaussian Markov chain is monotone, and for each $\xi$ has a unique initialization $c_{\xi}$ such that in $O(1)$ time, it exits the unstable fixed point to the right (towards the ordered phase) with probability $1-\xi$, and to the left (towards the disordered phase) with probability $\xi$. This behavior transfers to the CM dynamics by the approximation, and by picking the initialization’s edge-probability $\lambda_{0}/n$ so that the initial giant component size has mean $\theta_{*}n+c_{\xi}\sqrt{n}+o(\sqrt{n})$.
@@ 1.4.1 Escaping the unstable fixed point separating phases
then the giant component, centered by $\theta_{*}n$ and divided by $\sqrt{n}$, tracks this process closely. I.e., if we denote by $\bar{L}_{t}=n^{-1/2}(|\mathcal{L}_{1}(X_{t})|-\theta_{*}n)$, by Theorem 3.19 there exists a coupling $\mathbf{P}$ such that for every $T=O(1)$,
@@
With (4) in hand, we translate exit probabilities for $\bar{Z}_{t}$ to the right and left of an interval $[-\gamma,\gamma]$ with $\gamma$ large, to exit probabilities on $\bar{L}_{t}$ with only $o(1)$ error. A subtle technical point is that, although the variances of $\bar{Z}_{t}-\bar{Z}_{t-1}$’s steps are time-dependent, they are notably neither $n$-dependent, nor $\bar{Z}_{t-1}$-dependent, so this is a tractable and monotone chain on $\mathbb{R}$.
@@ 1.4.2 Mixing within a phase away from the fixed-point
From the above we deduce that for a well-chosen initialization parameter $\lambda_{0}$, after $O(1)$ many steps, $|\mathcal{L}_{1}(X_{t})|$ has exited $[\theta_{*}n-\gamma\sqrt{n},\theta_{*}n+\gamma\sqrt{n}]$ to the right with probability $1-\xi+o(1)$ and to the left with probability $\xi-o(1)$. The next step is to show that if it exited to the right, with probability $1-O(\gamma^{-2})$, in a further $O(\log n)$ steps, it quasi-equilibrates from there to the ordered phase (the random-cluster measure conditioned on having a giant), and if it exited to the left similarly to the disordered phase. Without loss of generality, let us discuss the first setting of exiting to the right.
@@
    initially, the drift away from the saddle for $|\mathcal{L}_{1}(X_{t})|$ is proportional to its distance to $\theta_{*}n$, and this can be used to get it from $\theta_{*}n+\gamma\sqrt{n}$ to $\theta_{*}n+\Omega(n)$ in $O(\log n)$ steps, except with probability $1-O(\gamma^{-2})$ (which covers the possibility it goes back to the saddle in which case quasi-equilibration fails.
@@
    any two configurations having $|\mathcal{L}_{1}(X_{t})-\theta_{\mathsf{r}}n|=O(\sqrt{n})$ and a few other easy to maintain properties can be coupled with probability $1-\varepsilon$ in $O_{\varepsilon}(1)$ further steps; since the ordered phase measure also satisfies these properties, we can couple $X_{t}$ to a sample from the ordered phase with high probability.
@@ 1.4.3 The Potts Glauber dynamics proof
Some notable further challenges arise in the low-temperature regime of $\beta>\beta_{\mathsf{s}}$ that are not present in the CM analysis, due to the fact that the saddle at $(1/q,...,1/q)$ proportions vector is unstable in all $q$ directions simultaneously, while when $\beta<\beta_{\mathsf{s}}$, the saddles we encounter are stable in all but one direction. To handle these difficulties, we obtain a better quantitative understanding of the $q$-dimensional drift landscape (for general $q$), and use refined approximations of the behavior of the Glauber dynamics’ proportions vector away from its fixed points by a $q$-dimensional gradient dynamical system. For the reader only interested in the Potts Glauber dynamics, this is found in Section 4, and its presentation is pretty much self-contained.
@@ Acknowledgments
The research of R.G. is supported in part by NSF grant DMS-2246780. The research of A.B. and X.Z. is supported in part by NSF grant CCF-2143762.
@@ 2 Preliminaries
Let us begin with some discussion of the notation used throughout the paper. We will always think of $\beta,q$ fixed, and therefore may drop their dependencies from subscripts. All our results should be thought of as applying for $n$ sufficiently large. When we use big-$O$, little-$o$, and $\Theta$ notation, the implicit constants may depend on $\beta,q$ but not on $n$, and also not on the threshold $\gamma$ which will appear frequently for the window around the saddle point we are zooming in around.
@@ 2.1 Structure of the mean-field Potts and random-cluster equilibrium measures
$\xi:=\frac{1}{1+\xi^{\prime}},\qquad\text{where}\qquad\xi^{\prime}:=\frac{1}{q-1}\Big{(}\frac{q-\beta_{\mathsf{c}}/(q-1)}{q-\beta_{\mathsf{c}}}\Big{)}^{(2-q)/2}\exp{\Big{(}\frac{\beta_{\mathsf{c}}^{2}(q-2)(q^{2}-4q+2)}{4q(q-1)^{2}}\Big{)}}.$ (6)
@@
  • If $\beta<\beta_{\mathsf{c}}$, with probability $1-o(1)$, it has
@@
  • If $\beta>\beta_{\mathsf{c}}$, with probability $1-o(1)$, it has
@@
  • If $\beta<\beta_{\mathsf{c}}$, with probability $1-o(1)$, it has
@@
  • If $\beta>\beta_{\mathsf{c}}$, with probability $1-o(1)$, (up to a permutation of the $q$ spins) it has
@@ 2.2 Random graph preliminaries
Due to the percolation step in the definition of the CM dynamics, precise random graph estimates are essential to careful understanding of the CM dynamics. We begin with central limit and local limit theorems for the size of the giant in a super-critical Erdős–Renyi random graph.
@@ 2.2.1 Limit theorems for the size of the giant component
For $\lambda>1$, the expected size of the giant component in a $G(n,\lambda/n)$ random graph is roughly $\alpha(\lambda)n$ where
$\alpha(\lambda):=\max\{x>0:e^{-\lambda x}=1-x\}\,.$ (7)
@@
[…]
@@ 2.2.2 Discrete duality
By a standard calculation, if $|U|=\alpha(\lambda)n+o(n)$, then $\frac{n-|U|}{n}\lambda<1$ is uniformly bounded away from $1$, so that the resulting graph is uniformly sub-critical and the probability of $\mathcal{E}_{|U|}$ is exponentially small in $n$. The result follows from Observation 2.5. ∎
@@ 2.2.3 Refined random graph statistics
$\mathbb{E}[R_{2}(G)]=\frac{n}{1-\lambda}\cdot\Big{(}1+O(n^{-1}(1-\lambda)^{-3})\Big{)}\,,\qquad\mathbb{E}[R_{3}(G)]=\frac{n}{(1-\lambda)^{3}}\cdot\Big{(}1+O(n^{-1}(1-\lambda)^{-3})\Big{)}\,,$ (12)
@@ 2.3 Drift of the giant component in the CM dynamics
We end this preliminaries section by describing properties of the drift function for the size of the giant component in the CM dynamics when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$. This function will govern the expected change to the (fractional) size of the giant component in the CM dynamics on the event that the giant is activated. To start, for $\theta\in[0,1]$, let
@@
so that $\phi(\theta)$ is the expected fractional size of the giant component from a configuration with fractional size $\theta$. Define the drift function
@@
Note that $f(\theta)$ is well-defined if $\theta\in(\theta_{\mathsf{s}},1]$, where $\theta_{\mathsf{s}}=\inf\{\theta:\beta k_{a}(\theta)>1\}=\frac{q-\beta}{\beta(q-1)}$; below $\theta_{\mathsf{s}}$, the percolation step will typically be sub-critical and in one step the fractional size of the giant would become $o(1)$, so we can formally set $\alpha(\theta)=0$ for $\theta<1$.

### v2

@@ 1.2 Initializations at other temperatures and relations to simulated annealing
Another central purpose of understanding mixing times of high-entropy initializations for Markov chains is that it is closely related to understanding simulated annealing schemes. Simulated annealing, introduced in [38] (and variants like simulated tempering [44]) lowers the temperature over the run of the Markov chain, so that when run at low temperatures, one is effectively initializing from the stationary measure at a nearby, but a higher-temperature (and therefore higher-entropy) distribution. This is one of the most commonly implemented modifications to vanilla MCMC to make it better able to sample from multimodal low-temperature landscapes. Simulated annealing and simulated tempering have been analyzed in the context of mean-field Potts dynamics by e.g., [1, 52], and it is known that the discontinuous phase transition at $\beta_{\mathsf{c}}$ presents a serious obstruction to fast mixing from such schemes, for instance because the high-temperature initialization is metastable when $\beta\in(\beta_{\mathsf{c}},\beta_{\mathsf{s}})$.
Still, our results lend insight into the approach by precisely classifying the set of temperatures $\beta_{0}$ for which initialization from $\mu_{\beta_{0}}$ or $\pi_{\beta_{0}}$ would be fast for the CM and Potts Glauber dynamics.
Theorem 1.3.
For every $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$, there exists explicit $b_{*}(\beta,q)\geq\beta_{\mathsf{c}}(q)$ such that:
  1. 1.
    If $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$, the CM dynamics initialized from $\mu_{\beta_{0}}$ for $\beta_{0}<b_{*}(\beta,q)$ will mix in $O(\log n)$ steps, but will take at least $\exp(\Omega(n))$ steps to attain $o(1)$ TV-distance for any $\beta_{0}>b_{*}(\beta,q)$.
  2. 2.
    If $\beta=\beta_{\mathsf{c}}$, the CM dynamics initialized from $\mu_{\beta_{0}}$ will take at least $\exp(\Omega(n))$ steps to attain $o(1)$ TV-distance for any $\beta_{0}\neq\beta_{\mathsf{c}}$.
  3. 3.
    If $\beta\in(\beta_{\mathsf{c}},\beta_{\mathsf{s}})$, the CM dynamics initialized from $\mu_{\beta_{0}}$ mixes in $O(\log n)$ steps for every $\beta_{0}>b_{*}(\beta,q)$ but will take at least $\exp(\Omega(n))$ steps to attain $o(1)$ TV-distance for any $\beta_{0}<b_{*}(\beta,q)$.
Theorem 1.4.
For every $\beta\geq\beta_{\mathsf{u}}$, there exists explicit $d_{*}(\beta,q)\geq\beta_{\mathsf{c}}(q)$ such that:
  1. 1.
    If $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$, the Potts Glauber dynamics initialized from $\pi_{\beta_{0}}$ for $\beta_{0}<d_{*}(\beta,q)$ will mix in $O(n\log n)$ steps, but take at least $\exp(\Omega(n))$ steps to attain $o(1)$ TV-distance for any $\beta_{0}>d_{*}(\beta,q)$.
  2. 2.
    If $\beta=\beta_{\mathsf{c}}$, the Potts Glauber dynamics initialized from $\pi_{\beta_{0}}$ will take at least $\exp(\Omega(n))$ steps to attain any $o(1)$ TV-distance for every $\beta_{0}\neq\beta_{\mathsf{c}}$.
  3. 3.
    If $\beta\in(\beta_{\mathsf{c}},\beta_{\mathsf{s}})$, the Potts Glauber dynamics initialized from $\pi_{\beta_{0}}$ mixes in $O(n\log n)$ steps for every $\beta_{0}>d_{*}(\beta,q)$ but will take at least $\exp(\Omega(n))$ steps to attain $o(1)$ TV-distance for any $\beta_{0}<d_{*}(\beta,q)$.
  4. 4.
    If $\beta>\beta_{\mathsf{s}}$, the Potts Glauber dynamics initialized from $\pi_{\beta_{0}}$ mixes in $O(n\log n)$ steps for every $\beta_{0}\geq 0$.
@@ 1.3 Proof outlines
In what follows, let $f(\theta)=f_{\beta,q}(\theta)$ be the expected drift for the size of the giant component in one step of CM dynamics. When $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$, let $\theta_{*}$ be its unstable fixed point separating the two stable fixed points of 0 (the disordered phase) and $\theta_{\mathsf{r}}$ (the ordered phase); see Figure 3 and Section 2.3 for the precise definitions. For intuition, when comparing to Figure 1, the drift is roughly the derivative of the log-probability landscape: a giant component of size $\theta_{*}n$ corresponds to the saddle point separating the modes whose minima are at no giant component, and at a giant component of size $\theta_{\mathsf{r}}n$.
@@
    Escape from a local neighborhood of the unstable fixed point: We show that in an $O(\sqrt{n})$ window around the unstable fixed point $\theta_{*}n$, the evolution of the size of its giant component, though not literally a Markov process, is well-approximated by a 1-dimensional Gaussian Markov chain (so long as the dynamics maintains a certain typical structure).
    The 1-dimensional Gaussian Markov chain is monotone in the initialization, and for each $\xi$ has a unique initialization $c_{\xi}$ such that in $O(1)$ time, it exits the unstable fixed point to the right (towards the ordered phase) with probability $1-\xi$, and to the left (towards the disordered phase) with probability $\xi$. This behavior transfers to the CM dynamics by the approximation, and by picking the initialization’s edge-probability $\lambda_{0}/n$ so that the initial giant component size has mean $\theta_{*}n+c_{\xi}\sqrt{n}+o(\sqrt{n})$.
@@ 1.3.1 Escaping the unstable fixed point separating phases
then the giant component, centered by $\theta_{*}n$ and divided by $\sqrt{n}$, tracks this process closely. I.e., if we let $\bar{L}_{t}=n^{-1/2}(|\mathcal{L}_{1}(X_{t})|-\theta_{*}n)$, then by Theorem 3.19 there exists a coupling $\mathbf{P}$ such that for every $T=O(1)$,
@@
With (4) in hand, we translate exit probabilities for $\bar{Z}_{t}$ to the right and left of an interval $\gamma,\gamma]$ with large $\gamma=O(1)$, to exit probabilities on $\bar{L}_{t}$ with only $o(1)$ error. A subtle technical point is that, although the variances of $\bar{Z}_{t}-\bar{Z}_{t-1}$’s steps are time-dependent, they are notably neither $n$-dependent, nor $\bar{Z}_{t-1}$-dependent, so this is a tractable and monotone chain on $\mathbb{R}$.
@@ 1.3.2 Mixing within a phase away from the fixed-point
From the above we deduce that for a well-chosen initialization parameter $\lambda_{0}$, after $O(1)$ many steps (depending on $\gamma$), $|\mathcal{L}_{1}(X_{t})|$ has exited $\theta_{*}n-\gamma\sqrt{n},\theta_{*}n+\gamma\sqrt{n}]$ to the right with probability $1-\xi+o(1)$ and to the left with probability $\xi-o(1)$. The next step is to show that if it exited to the right, with probability $1-O(\gamma^{-2})$, in a further $O(\log n)$ steps, it quasi-equilibrates from there to the ordered phase (the random-cluster measure conditioned on having a giant), and if it exited to the left similarly to the disordered phase. Without loss of generality, let us discuss the first setting of exiting to the right.
@@
    initially, the drift away from the saddle for $|\mathcal{L}_{1}(X_{t})|$ is proportional to its distance to $\theta_{*}n$, and this can be used to get it from $\theta_{*}n+\gamma\sqrt{n}$ to $\theta_{*}n+\Omega(n)$ in $O(\log n)$ steps, except with probability $1-O(\gamma^{-2})$, which covers the possibility it goes back to the saddle in which case quasi-equilibration fails.
@@
    any two configurations having $|\mathcal{L}_{1}(X_{t})-\theta_{\mathsf{r}}n|=O(\sqrt{n})$ and a few other easy-to-maintain properties can be coupled with probability $1-\varepsilon$ in $O_{\varepsilon}(\log n)$ further steps; since the ordered phase measure also satisfies these properties, we can couple $X_{t}$ to a sample from the ordered phase with high probability.
@@ 1.3.3 The Potts Glauber dynamics proof
Some notable further challenges arise in the low-temperature regime of $\beta>\beta_{\mathsf{s}}$ that are not present in the CM analysis, due to the fact that the saddle at (1/q,,1/q)(1/q,...,1/q) proportions vector is unstable in all $q$ directions simultaneously, while when $\beta<\beta_{\mathsf{s}}$, the saddles we encounter are stable in all but one direction. To handle these difficulties, we obtain a better quantitative understanding of the $q$-dimensional drift landscape (for general $q$), and use refined approximations of the behavior of the Glauber dynamics’ proportions vector away from its fixed points by a $q$-dimensional gradient dynamical system. For the reader only interested in the Potts Glauber dynamics, this is found in Section 4, and its presentation is effectively self-contained.
@@ Acknowledgments
The authors thank anonymous referees for their careful reading and helpful comments. The research of R.G. is supported in part by NSF grant DMS-2246780. The research of A.B. and X.Z. is supported in part by NSF grant CCF-2143762.
@@ 2 Preliminaries
Let us begin with some discussion of the notation used throughout the paper. We will always think of $\beta,q$ fixed, and therefore may drop their dependencies from subscripts. All our results should be thought of as applying for $n$ sufficiently large. When we use big-$O$, little-$o$, and $\Theta$ notation, the implicit constants may depend on $\beta,q$ but not on $n$. The threshold $\gamma$ will ultimately be taken to be a sufficiently large constant depending on the $\varepsilon$ total variation distance to the stationary measure we are aiming for. Therefore, we mostly treat $\gamma=O(1)$ except when it is important to ensure that large enough $\gamma$ suffices to achieve $\varepsilon$ total variation distance. We use the notation $\tilde{O}$ to suppress logarithmic factors.
@@ 2.1 Structure of the mean-field Potts and random-cluster equilibrium measures
$\xi$ $:=\frac{1}{1+\xi^{\prime}},\qquad\text{where}$ (6)
$\xi^{\prime}$ $:=\frac{1}{q-1}\Big{(}\frac{q-\beta_{\mathsf{c}}/(q-1)}{q-\beta_{\mathsf{c}}}\Big{)}^{(2-q)/2}\exp{\Big{(}\frac{\beta_{\mathsf{c}}^{2}(q-2)(q^{2}-4q+2)}{4q(q-1)^{2}}\Big{)}}.$
@@
  • If $\beta<\beta_{\mathsf{c}}$, with probability 1o(1)1-o(1), it has
@@
  • If $\beta>\beta_{\mathsf{c}}$, with probability 1o(1)1-o(1), it has
@@
  • If $\beta<\beta_{\mathsf{c}}$, with probability 1o(1)1-o(1), it has
@@
  • If $\beta>\beta_{\mathsf{c}}$, with probability 1o(1)1-o(1), (up to a permutation of the $q$ spins) it has
@@ 2.2 Random graph preliminaries
Due to the percolation step in the definition of the CM dynamics, precise random graph estimates are essential to careful understanding of the CM dynamics. We begin with central limit and local limit theorems for the size of the giant in a super-critical Erdős–Renyi random graph. (Recall that a $G(n,\lambda/n)$ random graph is said to be sub-critical when $\lambda<1$ and is called super-critical when $\lambda>1$.)
@@ 2.2.1 Limit theorems for the size of the giant component
The expected size of the giant component in a $G(n,\lambda/n)$ random graph is roughly $\alpha(\lambda)n$ where
$\alpha(\lambda):=\begin{cases}\max\{x>0:e^{-\lambda x}=1-x\}&\lambda>1\\ 0&\lambda\leq 1\end{cases}\,.$ (7)
@@
$\frac{|\mathcal{L}_{1}(G)|-\alpha(\lambda)n}{\sigma(\lambda)\sqrt{n}}\xrightarrow[D]{}\mathcal{N}(0,1),$ (11)
where $\xrightarrow[D$]{} denotes convergence in distribution, and $\mathcal{N}(0,1)$ is a standard normal.
@@ 2.2.2 Discrete duality
By a standard calculation, if $|U|=\alpha(\lambda)n+o(n)$, then $\frac{n-|U|}{n}\lambda<1$ is uniformly bounded away from $1$, so that the resulting graph is uniformly sub-critical and the probability of $\mathcal{E}_{|U|}$ is $1-\exp(-\Omega(n))$. The result follows from Observation 2.5. ∎
@@ 2.2.3 Refined random graph statistics
$\mathbb{E}[R_{2}(G)]$ $=\frac{n}{1-\lambda}\cdot\Big{(}1+O(n^{-1}(1-\lambda)^{-3})\Big{)}\,$, (12)
$\mathbb{E}[R_{3}(G)]$ $=\frac{n}{(1-\lambda)^{3}}\cdot\Big{(}1+O(n^{-1}(1-\lambda)^{-3})\Big{)}\,$,
@@ 2.3 Drift of the giant component in the CM dynamics
We end this preliminaries section by describing properties of the drift function for the size of the giant component in the CM dynamics when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$. This function will govern the expected change to the (fractional) size of the giant component in the CM dynamics on the event that the giant is activated. To start, for $\theta\in[0,1]$, let $k_{\mathsf{a}}(\theta)$ be the expected fraction of activated vertices if a giant of fractional size $\theta$ is activated, i.e.,
@@
so that $\phi(\theta)$ is the expected fractional size of the giant component from a configuration with a giant of fractional size $\theta$. Define the drift function
@@
Note that $\alpha(\lambda)$ represents the expected fraction of the giant component, as characterized in (7). When $\lambda=\beta k_{\mathsf{a}}(\theta)$, the regime $\lambda>1$ corresponds to $\theta\in(\theta_{\mathsf{s}},1]$, where $\theta_{\mathsf{s}}=\inf\{\theta:\beta k_{a}(\theta)>1\}=\frac{q-\beta}{\beta(q-1)}$.